tax:

Algebra 2 - Question 5/5

tax:

\[\frac{ 2 }{ x-3 } + \frac{ 5 }{ x + 1 } = 1\] There's the equation

tax:

\[x=2, x=5\] Those are my final two solutions for solving for x

tax:

@unknownnnnnn

unknownnnnnn:

yes right too

tax:

@unknownnnnnn wrote:
yes right too
Alright thanks!

Daze01:

bruh..stop posting question for no reason on here when i know u can just AI it.

unknownnnnnn:

@tax wrote:
@unknownnnnnn wrote:
yes right too
Alright thanks!
your welcome

yhyhR:

alr that's hard